OpenStudy (anonymous):

EASY QUESTION!!! Medal to the best!!! What is the new volume of a rectangular prism that is originally eighteen cubic meters and is then resized using a scale factor of 2 on each side length? nine cubic meters thirty-six cubic meters seventy-two cubic meters one hundred forty-four cubic meters

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

now, increase using a scale factor of 2, 

OpenStudy (solomonzelman):

So the formula for the first one is \[X \times Y \times Z\] and we know that it's equal to18. for the new prism, \[X \times 2 \times Y \times 2 \times Z \times 2\] this is same as\[X \times Y \times Z \times 2 \times 2 \times 2\] which is same as\[X \times Y \times Z \times 8\] we know that \[X \times Y \times Z =18\]so, \[18 \times 8\] and \[18 \times 8=144\] so the final answer is \[144m^2\] Or D.